# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 2 The Lee-Wick model
where $m_{N}=939$ MeV is the nucleon mass, $g$ is the strong interaction coupling constant, and $U(\phi)$ is the scalar field potential. The scalar field $\phi$ will be treated on the mean-field level. Then the nucleon spectrum is $\sqrt{m_{N}^{*2}+p^{2}}$ with the effective nucleon mass $m_{N}^{*}=m_{N}-g\phi$. While in the vacuum $\phi=0$, Lee and Wick suggested that in the nucleon medium at some critical density the effective nucleon mass $m_{N}^{*}=m_{N}-g\phi$ can reach zero. It would occur for $\phi_{c}=m_{N}/g$. The energy density gain at this transition can be estimated as
@@
$n_{c}=n_{c,0}\Big(1+\frac{3p_{{\rm F},c}}{4m_{N}}+\frac{9p_{{\rm F},c}^{2}}{20m_{N}^{2}}+O(p_{{\rm F},c}^{3}/m_{N}^{3})\Big),\quad n_{c,0}=\frac{U(\phi_{c})}{m_{N}}=\nu\frac{p^{3}_{{\rm F},c}}{3\pi^{2}}$ (3)
@@
Using typical values for the strong interaction constant $g\sim 10$ and the mass of the scalar meson, $m_{\phi}\sim m_{\sigma}\sim 700$ MeV, we obtain $n_{c}\sim 2n_{0}$. The energy gain per nucleon for $n>n_{c}$ and for $(n-n_{c})/n_{c}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1$ can be estimated as $\mathcal{E}=\Delta E/n\simeq-(1-n_{c}/n)m_{N}$ . These rough estimates show to a possibility that the ordinary nuclei may be exist in a metastable state and after a while they may undergo the first-order phase transition to the superdense ground state. The energy of each nucleon inside the anomalous nucleus is $p$, and outside $\sqrt{m_{N}^{2}+p^{2}}$. Thus, there is a potential barrier holding the nucleons in the given abnormal state.
@@
$U(\phi)=\frac{1}{2}\mu^{2}\phi^{2}-\frac{1}{\sqrt{2}}\mu\sqrt{\lambda}(1-\delta)\phi^{3}+\frac{1}{4}\lambda\phi^{4}\,,\quad\mu,\lambda>0\,$, (5)
@@
$\phi_{2}=\frac{\phi_{0}}{4}(3(1-\delta)+\sqrt{9(1-\delta)^{2}-8})=\phi_{0}(1-3\delta+O(\delta)),\quad\mbox{where}\quad\phi_{0}=\sqrt{2}\frac{\mu}{\sqrt{\lambda}}.$ (6)
@@
and plot $\Delta E(\phi,n)=E(\phi,n)-E(0,n)$ as a function of $\phi$ for various values of the nucleon density. In Fig. 3 we present the results for the set of parameters
@@
Figure 2: Difference in the energy density (10) due to the development of the scalar field as function of $\phi$ for various nucleon densities ($\nu=2$) calculated with the model parameters (11).
Figure 3: Effective nucleon mass as a function of the nucleon density, $\nu=2$. Solid line is calculated for the Lee-Wick model with parameters (11). Dashed line is drawn for $\delta=0.02$. Dash-dotted line corresponds to the RMF model (15) with parameters (16).
Figure 4: Binding energy per nucleon (12) for the nuclear matter ($\nu=2$) as a function of nucleon density. Difference in the energy density (10) due to the development of the scalar field is calculated for parameters (11). Solid line shows the result of the Lee-Wick model. Dashed lines include nucleon-nucleon interaction term (13). RMF curve shows RMF calculation with parameters (16).
The same year when Lee and Wick proposed their model, Walecka formulated a relativistic mean-field model (RMF) Walecka (1974); Chin and Walecka (1974). Here one should also mention previous works by Johnson and Teller, and by Dürr Johnson and Teller (1955); Duerr (1956), where the nuclear saturation arises due to the difference between scalar and vector nucleon densities. The Walecka model includes nucleons and mean fields of $\sigma$ and vector $\omega$ mesons allowing the $\sigma$ field dependence of the nucleon mass. Coupling constants of $\sigma$ and vector $\omega$ mesons with nucleons were fitted to describe binding energy and saturation density of nuclear matter. Incompressibility at nuclear saturation proved to be unrealistically high and effective nucleon mass low. To improve the fit Boguta included possible scalar field self-interactions depending on two extra fitting parameters Boguta and Bodmer (1977). Also $\rho$ meson mean field was included. These changes made the RMF models more flexible, see Serot and Walecka (1997). If in Eq. (10) we make replacements
$\phi\to\sigma\,,\quad\mu\to m_{\sigma}\,,\quad g\to g_{\sigma}\,,\quad\lambda\to cg_{\sigma}^{4}\,\quad\delta\to 1+\frac{\sqrt{2}}{3\sqrt{c}m_{\sigma}}bg_{\sigma}m_{N}\,$, (14)
and add the repulsive contribution due to the $\omega$ meson exchange with the mass $m_{\omega}$ and the coupling constant ($g_{\omega}$), we obtain the RMF energy density
$E(\sigma,n)=\frac{m_{\sigma}^{2}\sigma^{2}}{2}+\frac{b}{3}m_{N}(g_{\sigma}\sigma)^{3}+\frac{c}{4}(g_{\sigma}\sigma)^{4}+\nu\intop_{0}^{p_{\rm F}}\frac{{\rm d}pp^{2}}{\pi^{2}}\sqrt{(m_{N}-g_{\sigma}\sigma)^{2}+p^{2}}+\frac{g_{\omega}^{2}n^{2}}{2m_{\omega}^{2}}\,.$ (15)
@@
As example, taking the nucleon binding energy $\mathcal{E}_{0}=16$ MeV, compressibility modulus $K=275$ MeV and the effective nucleon mass $m_{N}^{*}(n_{0})=0.805$ at the saturation density $n_{0}$, as we did in Ref. Kolomeitsev and Voskresensky (2005), we obtain parameters
$\frac{g_{\sigma}^{2}m_{N}^{2}}{m_{\sigma}^{2}}=184.356\,,\quad\frac{g_{\omega}^{2}m_{N}^{2}}{m_{\omega}^{2}}=87.5996\,,\quad b=5.53871\times 10^{-3}\,,\quad c=2.29759\times 10^{-2}\,.$ (16)
@@
The density dependence of the effective nucleon mass within the RMF model is shown by the dash-dotted line in Fig. 3. The mass decreases smoothly with the density increase without any jumps in difference to the Lee-Wick model. The binding energy per nucleon (12) for the nuclear matter ($\nu=2$) as a function of nucleon density is depicted in Fig. 4, solid line for Lee-Wick model. Long and short dashed lines include the nucleon-nucleon interaction term (13). The dash-dotted line shows the RMF calculation with parameters (16). The Lee-Wick abnormal state does not appear in the latter model.
Various modifications of the original Walecka model including $\rho$ mesons, a non-linear $\sigma$ field dependent non-linear potential, possible $\omega-\sigma$ interaction, etc, are widely and rather successfully used to describe equation of state of the hadron matter in compact stars, heavy-ion collisions and finite nuclei, cf. Serot and Walecka (1997); Geng et al. (2005); Voskresensky (1993) and references therein. In spite of all these goodies, for densities essentially above $n_{0}$ these models are phenomenological extrapolations. With parameters which were fitted to describe the nuclear system at $n=n_{0}$, at larger densities in these models the Lee-Wick abnormal state does not appear. However information about chiral symmetry is lost in these models, whereas the Lee and Wick model can be treated as a chiral symmetry based version of the RMF model. Pions are not included in both the Lee–Wick and RMF approaches, since they do not contribute on the mean-field level. Although not permitting to describe properties of the ordinary atomic nuclei and nuclear matter at $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{0}$, the Lee-Wick model demonstrates principle possibility of existence of the new stable (or metastable) abnormal nuclear state at $n$ significantly above $n_{0}$. Note that one of ways to study the possibility of the Lee-Wick second minimum in $V(\sigma)$ within the relativistic mean-field approach to the equation of state of hadronic matter could be to employ within the latter approach the idea of a partial conservation of the chiral symmetry. It was realized in a series of works Kolomeitsev and Voskresensky (2005); Maslov et al. (2015, 2016); Kolomeitsev et al. (2017), where it was assumed that not only baryon but all hadron (baryon, $\omega$, $\rho$ and the $\sigma$ field) mass-terms decrease with increasing $\sigma$ field. Such assumption have permitted to successfully resolve several problems (associated with existence of very massive neutron stars, absence of the efficient direct Urca neutrino reaction in most neutron stars, so called hyperon and $\Delta$ isobar puzzles). For simplicity the parameters were chosen, which do not allow for the new minimum. Thus question about possibility of the Lee-Wick state within such kind of approach still remains open.
3 Supercharged nuclei, nuclearites and nuclei–stars
3.1 Idea of supercharged nuclei in a simplified treatment
For a moment, let us assume that negatively charged light scalar bosons (or pseudoscalar "light pions") exist with a mass of $m_{b}$. Could these bosons bind supercharged nuclei, as was suggested in Voskresensky (1977), see also review Migdal et al. (1990)? The energy of the nucleus with approximately constant density $n\simeq n_{0}$ at the proton number $Z\simeq A/2$ is given by
@@ 3.1 Idea of supercharged nuclei in a simplified treatment
$+\int\left(|\nabla\phi|^{2}-V^{2}|\phi|^{2}+m_{b}^{2}+{\textstyle\frac{1}{2}}\lambda|\phi|^{2})|\phi|^{2}\right){\rm d}^{3}x\,,\quad\mathcal{E}_{0}=16\,{\rm MeV}.$ (17)
Here the first term is the volume binding energy, the second term produces the surface energy, which is suppressed as $1/A^{1/3}$ compared to the first term and will be dropped when the nucleus atomic number $A$ is large as we assume. The proton density is $n_{p}\simeq n_{0}/2$. The third and fourth terms describe Coulomb interactions of protons and the electric field energy For ordinary atomic nuclei these terms produce the Coulomb energy $3Z^{2}e^{2}/5R$, where the nucleus radius is $R\simeq r_{0}A^{1/3}=(r_{0}/2^{1/3})Z^{1/3}$ and $r_{0}\simeq 1.2$ fm. The last term is the contribution of the classical charged boson field, the coupling of the pion-pion self-interaction $\lambda\sim 1$. With an increase in charge $Z\sim A/2$, the Coulomb energy repulsion always begins to exceed the volume energy attraction, so for the existence of extended systems with $Z\sim A/2>>1/e^{3}$ a charge must be ar least partially compensated. When $A$ and $Z$ grow, the electrical potential well deepens. When it becomes $V\leq-m_{b}$, the charged light bosons may appear in the reactions $n\to p+b^{-}$ and compensate the charge of the protons. As the result, in a nucleus with $A>>1/e^{3}$, the charge-neutral interior is formed Migdal et al. (1977) surrounded by a thin surface layer, where the strength of the electrical field is $E\sim 10^{2}m_{e}^{2}\sim 10^{18}$ V/cm with $m_{e}=0.5$ MeV being the electron mass.
@@
From these equations in case of a spherical nucleus with large $A$ we get uniform solutions $V=-m_{b}$ and $|\phi|^{2}=n_{p}/(2m_{b})$ valid in the interior, except for a boundary layer of the depth determined by the typical Debye screening length, being much shorter than the nucleus size. Here we put $\lambda\to 0$ to simplify consideration. Hence after the formation of the charged boson condensate the potential well stops growing keeping limiting depth of $V=-m_{b}$. Thus the volume part of the energy reads now
@@
We conclude that if negatively charged scalar bosons with a mass of $m_{b}<2\mathcal{E}_{0}\simeq(30\mbox{--}32)$ MeV existed in nature, then stable supercharged nuclei and nuclearites of an arbitrary size (with $A$ up to $10^{56}$ and $n\sim n_{0}$) would also exist, being glued by strong and electromagnetic forces, see Refs. Voskresensky (1977); Migdal et al. (1990).
These exotic objects have been given various names: anomalous nuclei, abnormal nuclei, nuclear drops, and nuclei-stars. Similar self-bound objects made of quark were also called — nuggets, nuclearites, or strangelets. For larger-$A$ systems, gravity comes into the game, and we deal with compact stars with nucleon–light-boson (e.g., light pion) component. For objects build with the Witten’s strange quark matter, we will use further on the names like strangelets and strange stars.
@@
3.2 Dibaryons and neutral light bosons

### v2

@@ II The Lee-Wick model
where $m_{N}=939$ MeV is the nucleon mass, $\psi$ is the nucleon bispinor field, $g$ is the strong interaction coupling constant, and $U(\phi)$ is the scalar field potential. The scalar field $\phi$ will be treated on the mean-field level. Then the nucleon spectrum is $\sqrt{m_{N}^{*2}+p^{2}}$ with the effective nucleon mass $m_{N}^{*}=m_{N}-g\phi$. While in the vacuum $\phi=0$, Lee and Wick suggested that in the nucleon medium at some critical density the effective nucleon mass $m_{N}^{*}(n)$ can reach zero. It would occur for $\phi_{c}=m_{N}/g$. The energy density gain at this transition can be estimated as
@@
$n_{c}$ $=n_{c,0}\Big(1+\frac{3p_{{\rm F},c}}{4m_{N}}+\frac{9p_{{\rm F},c}^{2}}{20m_{N}^{2}}+O(p_{{\rm F},c}^{3}/m_{N}^{3})\Big),\quad$
$n_{c,0}$ $=\frac{U(\phi_{c})}{m_{N}}=\nu\frac{p^{3}_{{\rm F},c}}{3\pi^{2}}\,.$ (3)
@@
Using typical values for the strong interaction constant $g\sim 10$ and the mass of the scalar meson, $m_{\phi}\sim m_{\sigma}\sim 700$ MeV, we obtain $n_{c}\sim 2n_{0}$. The energy gain per nucleon for $n>n_{c}$ and for $(n-n_{c})/n_{c}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1$ can be estimated as $\mathcal{E}=\Delta E/n\simeq-(1-n_{c}/n)m_{N}$ . These rough estimates show to a possibility that the ordinary nuclei may be exist in a metastable state and after a while they may undergo the first-order phase transition to the superdense ground state. The energy of each nucleon inside the abnormal nucleus is $p$, and outside $\sqrt{m_{N}^{2}+p^{2}}$. Thus, there is a potential barrier holding the nucleons in the given abnormal state.
@@
$U(\phi)=\frac{1}{2}\mu^{2}\phi^{2}-\frac{\mu\sqrt{\lambda}}{\sqrt{2}}(1-\delta)\phi^{3}+\frac{1}{4}\lambda\phi^{4}\,,\,\,\mu,\lambda>0\,$, (5)
@@
$\phi_{2}$ $=\frac{\phi_{0}}{4}(3(1-\delta)+\sqrt{9(1-\delta)^{2}-8})$
$=\phi_{0}(1-3\delta+O(\delta)),\,\,\mbox{where}\quad\phi_{0}=\sqrt{2}\frac{\mu}{\sqrt{\lambda}}.$ (6)
@@
and plot $\Delta E(\phi,n)=E(\phi,n)-E(0,n)$ as a function of $\phi$ for various values of the nucleon density. In Fig. 2 we present the results obtained for isospin-symmetric matter (ISM), $\nu=2$, and for the set of parameters
@@
Figure 2: Difference in the energy density (10) due to the development of the scalar field for ISM as function of $\phi$ for various nucleon densities calculated with the model parameters (11).
Figure 3: Effective nucleon mass for ISM as a function of the nucleon density. Solid line is calculated for the Lee-Wick model with parameters (11). Dashed line is drawn for $\delta=0.02$. Dash-dotted line corresponds to the RMF model (15) with parameters (16).
Figure 4: Binding energy per nucleon (12) for ISM as a function of nucleon density. Difference in the energy density (10) due to the development of the scalar field is calculated for parameters (11). Solid line shows the result of the Lee-Wick model. Dashed lines include nucleon-nucleon interaction term (13). The dash-dotted curve shows the RMF calculation with parameters (16).
The same year when Lee and Wick proposed their model, Walecka formulated a relativistic mean-field model (RMF) [295, 52]. Here one should also mention previous works by Johnson and Teller, and by Dürr [136, 74], where the nuclear saturation arises due to the difference between scalar and vector nucleon densities. The Walecka model includes interaction of nucleons with mean fields of the scalar $\sigma$ and vector $\omega$ mesons allowing the $\sigma$ field dependence of the nucleon mass. Coupling constants of $\sigma$ and vector $\omega$ mesons with nucleons were fitted to describe binding energy and saturation density of nuclear matter. Incompressibility at nuclear saturation proved to be unrealistically high and effective nucleon mass low. To improve the fit Boguta included possible scalar field self-interactions depending on two extra fitting parameters [35]. Also $\rho$ meson mean field was included. These changes made the RMF models more flexible, see [259]. Doing in Eq. (10) replacements
$\phi$ $\to\sigma\,,\quad\mu\to m_{\sigma}\,,\quad g\to g_{\sigma}\,$$\quad$
$\lambda$ $\to cg_{\sigma}^{4}\,\quad\delta\to 1+\frac{\sqrt{2}}{3\sqrt{c}m_{\sigma}}bg_{\sigma}m_{N}\,$, (14)
and adding the repulsive contribution due to the $\omega$ meson exchange with the mass $m_{\omega}$ and the coupling constant $g_{\omega}$, we obtain the RMF energy density
$E(\sigma,n)$ $=\frac{m_{\sigma}^{2}\sigma^{2}}{2}+\frac{b}{3}m_{N}(g_{\sigma}\sigma)^{3}+\frac{c}{4}(g_{\sigma}\sigma)^{4}$
$+\nu\intop_{0}^{p_{\rm F}}\frac{{\rm d}pp^{2}}{\pi^{2}}\sqrt{(m_{N}-g_{\sigma}\sigma)^{2}+p^{2}}+\frac{g_{\omega}^{2}n^{2}}{2m_{\omega}^{2}}\,.$ (15)
@@
As example, taking the parameter of binding energy per nucleon $\mathcal{E}_{0}=16$ MeV, compressibility modulus $K=275$ MeV and the effective nucleon mass $m_{N}^{*}(n_{0})=0.805$ at the saturation density $n_{0}$, as we did in Ref. [154], we obtain parameters
$\frac{g_{\sigma}^{2}m_{N}^{2}}{m_{\sigma}^{2}}=184.356\,,\quad\frac{g_{\omega}^{2}m_{N}^{2}}{m_{\omega}^{2}}=87.5996\,$$\quad$
$b=5.53871\times 10^{-3}\,,\quad c=2.29759\times 10^{-2}\,.$ (16)
@@
The density dependence of the effective nucleon mass of the ISM calculated within the RMF model is shown by the dash-dotted line in Fig. 3. The mass decreases smoothly with the density increase without any jumps in difference to the Lee-Wick model. The binding energy per nucleon (12) for the nuclear matter ($\nu=2$) as a function of nucleon density is depicted in Fig. 4, solid line for Lee-Wick model. Long and short dashed lines include the nucleon-nucleon interaction term (13). The dash-dotted line shows the RMF calculation with parameters (16). The Lee-Wick abnormal state does not appear in the latter model.
Various modifications of the original Walecka model including $\rho$ mesons, a non-linear $\sigma$ field dependent non-linear potential, possible $\omega-\sigma$ interaction, etc, are widely and rather successfully used to describe equation of state of the hadron matter in compact stars, heavy-ion collisions and finite nuclei, cf. [259, 97] and references therein. Usually, one assumes that there is no pseudoscalar mean-field of pions in the spin balanced system. However, including pionic excitations beyond the mean-field level improves the description of the equation of state of nuclear matter at finite temperatures, e.g., in application to heavy-ion collisions [284]. At low temperatures, which are of main our interest below, the contribution of pions to the equation of state is rather small.
In spite of the success in describing various phenomena at densities $\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}2\,n_{0}$, RMF models should be considered as merely phenomenological extrapolations at densities $\gg n_{0}$. In Fig. 5 we compare the potentials of the scalar field for the Lee-Wick model and the RMF model. Dots show the values of the dimensionless magnitude of the scalar field for five values of the nucleon density from $1\,n_{0}$ till $5\,n_{0}$ with the step $1\,n_{0}$. For parameters fitted to the properties of the nuclear system at $n=n_{0}$, the Lee-Wick abnormal state does not appear at larger densities in the RMF model. As seen in Fig. 5, no information about the possible restoration of the chiral symmetry is included in the standard RMF models, whereas a mechanism for a phase transition to a new state with a small nucleon mass is build in the Lee-Wick model by construction. Although the latter model cannot describe the properties of the ordinary atomic nuclei and nuclear matter at $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{0}$, it demonstrates the principle possibility of the existence of a new stable (or metastable) abnormal nuclear state at $n$ significantly above $n_{0}$.
Figure 5: Potentials of the scalar field for the Lee-Wick model (5) with parameters (11) and for the RMF model with parameters (16) as functions of the dimensionless scalar field. Dots show the values of the dimensionless magnitude of the scalar field for five values of the nucleon density from $1\,n_{0}$ till $5\,n_{0}$ with the step $1\,n_{0}$.
[…]
III Supercharged nuclei, nuclearites and nuclei–stars
In Ref. [214, 211] three types of abnormal nuclei with a $\pi$ condensate were discussed. These were named the “supercharged”, “neutron” and “superheavy” anomalous nuclei. Later, it became clear that this gradation is very conditional. What is important is that these hypothetical objects may differ in their baryon number, interior density, isospin content, and the type of the condensate ($\pi^{\pm}$ or $\pi^{0}$) that holds them together.
III.1 Supercharged nuclei in a nutshell
We consider a nuclear system with a rather large baryon number $A$, approximately constant density $n\sim n_{0}$ at the proton number $Z\approx A/2$ (in this respect the nucleus with a large $Z\gg 1/e^{2}=137$ is supercharged although for $Z\gg 1/e^{3}\sim 10^{3}$ the charge is compensated by electrons produced from vacuum, see below). To demonstrate the key idea, we assume for a moment that there exists a negatively charged light scalar bosons (or pseudoscalar ”light pions”) with a mass $m_{b}$. Below we will show that in realistic situations, pions, which mass is changed in nuclear matter and in external fields, can play the role of such a light boson. One may ask whether these light bosons can bind supercharged nuclei, see [282], and [210].
The energy of the nucleus with approximately constant density $n\simeq n_{0}$ at the proton number $Z\simeq A/2$ is given by
@@ III.1 Supercharged nuclei in a nutshell
$+\int\left(|\nabla\phi|^{2}-V^{2}|\phi|^{2}+m_{b}^{2}+{\textstyle\frac{1}{2}}\lambda|\phi|^{2})|\phi|^{2}\right){\rm d}^{3}x\,$$\quad$
$\mathcal{E}_{0}=16\,{\rm MeV}.$ (17)
Here the first term is the volume binding energy, the second term produces the surface energy, which is suppressed as $1/A^{1/3}$ compared to the first term and will be dropped when the nucleus atomic number $A$ is large as we for simplicity will assume. The proton density is $n_{p}\simeq n_{0}/2$. The third and fourth terms describe Coulomb interactions of protons with the electric field. For ordinary atomic nuclei these terms produce the Coulomb energy $3Z^{2}e^{2}/5R$, where the nucleus radius is $R\simeq r_{0}A^{1/3}=(r_{0}/2^{1/3})Z^{1/3}$ and $r_{0}\simeq 1.2$ fm. The last term in Eq. (17) is the contribution of the classical charged boson field, the coupling of the pion-pion self-interaction $\lambda\sim 1$. With an increase in charge $Z\sim A/2$, the Coulomb energy repulsion always begins to exceed the volume energy attraction. So for the existence of extended systems with $Z\sim A/2>>1/e^{3}$ the charge must be essentially compensated by electrons and $\mu^{-}$. Indeed when $A$ and $Z$ grow, the electrical potential well, $V$, deepens. When it becomes $V\leq-m_{b}$, the charged light bosons may appear in the reactions $n\to p+b^{-}$ and compensate the charge of the protons. As the result, in a nucleus with $Z>>1/e^{3}$, the charge-neutral interior is formed [209] surrounded by a thin (with thickness $\sim 10$ fm) surface layer, where the strength of the electrical field is $E\sim 10^{2}m_{e}^{2}\sim 10^{18}$ V/cm with $m_{e}=0.5$ MeV being the electron mass.
@@
Here we put $\lambda\to 0$ to simplify consideration. Hence From these equations in case of a spherical nucleus with large $A$ we get uniform solutions $V=-m_{b}$ and $|\phi|^{2}=n_{p}/(2m_{b})$ valid in the interior, except for a boundary layer of the depth determined by the typical Debye screening length, being much shorter than the nucleus size. Thus, after the formation of the charged boson condensate the potential well stops growing keeping limiting depth of $V=-m_{b}$.
The volume part of the energy reads now
@@
We conclude that if negatively charged scalar bosons with a mass of $m_{b}<2\mathcal{E}_{0}\simeq(30\mbox{--}32)$ MeV existed in nature, then stable supercharged nuclei and nuclearites of an arbitrary size (with $A$ up to $10^{56}$ and $n\sim n_{0}$) would also exist, being glued by strong and electromagnetic forces, see Refs. [282, 210]. For larger-$A$ systems, gravity comes into the game, and we deal with compact stars containing a nucleon–light-boson (e.g., light pion) component.
@@
In the literature the exotic objects have been given various names: anomalous nuclei, abnormal nuclei, nuclear drops, nuclearites, and nuclei-stars for heavy objects. Similar self-bound objects made of quarks were called — quark nuggets or strangelets, quark nuclearites, strange stars. For objects build with the Witten’s strange quark matter, we will use further on the names like strangelets, quark nuclearites and strange stars.
III.2 Dibaryons and neutral light bosons